OpenStudy (anonymous):

x-intercept and y-intercept of the equation 3x − 2y = 6

OpenStudy (shamim):

if u put x=0 in ur equation then u will get y interception

OpenStudy (shamim):

ur given equation is\[3x-2y=6\]

OpenStudy (shamim):

if i put x=0\[3 \times 0-2y=6\]

OpenStudy (shamim):

\[y=-3\]

OpenStudy (shamim):

its y interception

OpenStudy (anonymous):

x=2?

OpenStudy (dan815):

when y = 0 you can see what value you get for x that is the x intercept.. and when x = 0 you can see the y intercept

OpenStudy (shamim):

and if u put y=0 then u will get x interception

OpenStudy (anonymous):

Okay. Thanks.

OpenStudy (shamim):

welcome

OpenStudy (dan815):